OpenStudy (anonymous):

If tan x = .67, what is cot x

OpenStudy (anonymous):

\[\huge \color{green}{cotx = \frac{1}{tanx}}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Cot = 1/tan = 1/0.67